OpenStudy (anonymous):

simplify 5^2-3 times 2

OpenStudy (anonymous):

25-6 19

OpenStudy (angela210793):

is it 5^2-3*2?

OpenStudy (angela210793):

or (5^2-3)*2?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

19

OpenStudy (lalaly):

\[5^2-(3 \times 2) = 25-6 = 19\]

OpenStudy (angela210793):

5^2-(3*2)... he said

OpenStudy (anonymous):

Gahhh nvm